Managerial Accounting

Please describe the circumstances of the following case study and recommend a course of action. Explain your approach to the problem, perform relevant per unit calculations and analysis, and formulate a recommendation. Ensure your work and recommendation are thoroughly supported.

Case Study:

Papaya Partners is a distributor of papayas. They purchase papayas from individual growers and package them in 10-pound cartons for delivery to their various customers, generally supermarkets. Last month, they budgeted to sell $500,000 worth of cartons at a price of $25 each. Actual sales met a budget of $500,000 at $25 per carton.

Management has received cost information based on actual performance and needs to understand the drivers of the overall variance from the budget. They have asked you, as an analyst in their management accounting department, to calculate and explain the variances. The following data has been provided:


Budget
Cost of fruit @ 10 pounds per carton   $      200,000
Cost of packaging @ 1 pound per carton   $        10,000
Labor costs @ .5 hourse per carton   $        90,000
Total Cost   $      300,000


Actual 
Cost of fruit @ 10 pounds per carton   $      244,200
Cost of packaging @ .55 pound per carton   $        11,000
Labor costs @ .75 hourse per carton   $      150,000
Total Cost   $      405,200


Unfavorable variance   $105,200.00



Specifically, management needs to know the:


  • Standard cost per unit (carton)

  • Actual cost per unit

  • Direct materials price variances

  • Direct materials usage variances

  • Direct labor rate variance

  • Direct labor efficiency variance


In addition, they would like to understand how the variances are calculated and what caused them. They would also like a recommendation on what might be done to improve the variances.

For this assignment, compute all required amounts and explain how the computations were performed. Describe whom you would work with to determine the causes of the variances and hypothesize on what caused the variances. Based on your analysis, recommend actions that management could take to improve the variances.

Superior papers will:


  • Perform all per unit calculations correctly.

  • Articulate how the calculations were performed.

  • Assess the variances computed and evaluate the operational results (i.e., is performance better or worse than budgeted?).

  • Explain with whom you would work to identify the root causes of the variances.

  • Propose well-thought-out causes for each variance.

  • Conclude on which variances require management’s attention and recommend courses of action.
Answer #1

I am solving 4 sub parts of the question:

1. Standard cost per unit = Standard cost is the estimated cost per unit including material, overheads and labor.

Here standard cost for 5,00,000 cartons is $3,00,000

So standard cost per unit = Total budget cost / budgeted number of cartons = $ 3,00,000 / 5,00,000 = $ 0.60 per carton

2. Actual cost per unit = Total actual cost / actual number of cartons = $ 4,05,200 / 5,00,000 = $ 0.8104 per carton

3. Direct material price variance = It is the difference between actual amount spent on direct material and the amount that would have been spent if the material had been purchased at standard price

So direct material price variance = ($ 244200+$ 11000) - ($ 200000+$ 10000) = $ 45,200 (Unfavorable)

4. Direct material usage variance - It is the difference between actual quantity of material utilized during a period and the standard consumption of material for the level of output achieved

Direct material price variance = (Actual quantity - standard quantity) X Standard price

Actual quantity Standard Quantity
Fruits
Total cost 244200 200000
per carton cost 10 10
Quantity 24420 20000
Packaging
Total cost 11000 10000
per carton cost 0.55 1
Quantity 20000 10000
Direct material price variance
Fruit (24420 - 20000) * 10 = $ 44200 (U)
Packaging (20000 - 10000) *1 = $ 10000 (U)